Factor completely:
x^2 - 5x - 7x + 35 = x^2 - 12x + 35
2x^2 + 11x + 5
x^4y - 16y = x^4 - 15y
Thank you so much!

Found 2 solutions by jim_thompson5910, RAY100:
Answer by jim_thompson5910(13794)

I'm assuming that you want to factor




Looking at the expression , we can see that the first coefficient is , the second coefficient is , and the last term is .


Now multiply the first coefficient by the last term to get .


Now the question is: what two whole numbers multiply to (the previous product) and add to the second coefficient ?


To find these two numbers, we need to list all of the factors of (the previous product).


Factors of :
1,5,7,35
-1,-5,-7,-35


Note: list the negative of each factor. This will allow us to find all possible combinations.


These factors pair up and multiply to .
1*35
5*7
(-1)*(-35)
(-5)*(-7)

Now let's add up each pair of factors to see if one pair adds to the middle coefficient :


First NumberSecond NumberSum
1351+35=36
575+7=12
-1-35-1+(-35)=-36
-5-7-5+(-7)=-12



From the table, we can see that the two numbers and add to (the middle coefficient).


So the two numbers and both multiply to and add to


Now replace the middle term with . Remember, and add to . So this shows us that .


Replace the second term with .


Group the terms into two pairs.


Factor out the GCF from the first group.


Factor out from the second group. The goal of this step is to make the terms in the second parenthesis equal to the terms in the first parenthesis.


Combine like terms. Or factor out the common term

---------------------------------------------


Answer:


So factors to .


Note: you can check the answer by FOILing to get or by graphing the original expression and the answer (the two graphs should be identical).








# 2




Looking at the expression , we can see that the first coefficient is , the second coefficient is , and the last term is .


Now multiply the first coefficient by the last term to get .


Now the question is: what two whole numbers multiply to (the previous product) and add to the second coefficient ?


To find these two numbers, we need to list all of the factors of (the previous product).


Factors of :
1,2,5,10
-1,-2,-5,-10


Note: list the negative of each factor. This will allow us to find all possible combinations.


These factors pair up and multiply to .
1*10
2*5
(-1)*(-10)
(-2)*(-5)

Now let's add up each pair of factors to see if one pair adds to the middle coefficient :


First NumberSecond NumberSum
1101+10=11
252+5=7
-1-10-1+(-10)=-11
-2-5-2+(-5)=-7



From the table, we can see that the two numbers and add to (the middle coefficient).


So the two numbers and both multiply to and add to


Now replace the middle term with . Remember, and add to . So this shows us that .


Replace the second term with .


Group the terms into two pairs.


Factor out the GCF from the first group.


Factor out from the second group. The goal of this step is to make the terms in the second parenthesis equal to the terms in the first parenthesis.


Combine like terms. Or factor out the common term

---------------------------------------------


Answer:


So factors to .


Note: you can check the answer by FOILing to get or by graphing the original expression and the answer (the two graphs should be identical).








# 3

I'm assuming you want to factor


Start with the given expression


Factor out the GCF


Factor using the difference of squares formula


Factor using the difference of squares formula


So completely factors to

Answer by RAY100(1637)

(1) x^2 -12x +35
(x-5)(x-7)
.
(2) 2X^2 +11x +5
(2x+1)(x+5)
.
(3) x^4y -16y = x^4 -15y
x^4y -16y +15y -x^4 =0
x^4y -1y -x^4 =0
x^4 (y-1) -y =0, easy answer
.
or
x^4y -16y+15y -x^4 =0
x^4y -1y -x^4 =0
y(x^4 -1) -x^4 =0
y(x^2 +1)(x^2-1) -x^4 =0
y(x^2 +1)(x+1)(x-1) -x^4 =0,,,,better answer
.
checking
Let x=1, y=2,
original is (-1), answer is (-1) ,,,,ok